OpenStudy (anonymous):

HELP PLEASE!!! I WILL GIVE A MEDAL!!! Part A: Solve A = 9 over 2(x + 23) for x. (4 points) Part B: Determine the value of x when A = 108. (2 points) Part C: Solve –np – 90 > 30 for n. Show your work. (4 points)

OpenStudy (anonymous):

\[ A = \frac{9}{2(x+23)}\] So what's the first step? :)

OpenStudy (anonymous):

How could you start to get x by itself?

OpenStudy (anonymous):

This is for part A right?

OpenStudy (anonymous):

For C) - np - 90 > 30 "negative np minus 90 is greater than 30" we can add np to each side, giving us -90 > 30 + np "negative 90 is greater than 30 plus np" flip the inequality around, so it reads as np + 30 < -90 "np plus 30 is less than negative 90" Then subtract thirty and divide by p, making sure the final answer takes the form n < ........ "n is less than....."

OpenStudy (anonymous):

Yeah, that first equation was for part A

OpenStudy (anonymous):

Okay. Uh? well we have to find the variable that needs to be isolated which is "x". Dont we have to work bacwkards with PEMDAS? @AllTehMaffs

OpenStudy (anonymous):

Yeah. You can make your life much simpler though by making it looks like this then. \[ (A) = \frac{(9)}{(2)(x+23)}\] The parenthesis break things up into manageable chunks. You don't have any exponents greater than 1 so we're good there, so we can jump straight into multiplication and division. Remember that a term in a parenthesis can be kept together, se can multiply each side right off the bat by (x+23) \[(x+23)(A) = \frac{(9)}{(2)}\] What's a good next step then?

OpenStudy (anonymous):

okay um? sorry Im not very good at this so it may take me some time to figure out.

OpenStudy (anonymous):

:) no worries. Did that first step make sense?

OpenStudy (anonymous):

not really... why did you move (x + 23) from the right side to the left?

OpenStudy (anonymous):

I did that to get it out of the denominator. I didn't multiply through with the 2 while it was on the bottom because later in the problem you would then have to divide by 2. ie \[ A = \frac{9}{2(x+23)} \] \[ A = \frac{9}{2x + 46} \] \[ (2x + 46) A = 9 \] if you do it that way, the 2 in front of the x will have to get divided out anyway, so might as well keep them separated in the beginning :) Does that make sense? So yeah, the expression we're working with now is \[ (x+23) \cdot A = \frac{9}{2} \]

OpenStudy (anonymous):

oh.. Now I get it a little bit better. lol thankss

OpenStudy (anonymous):

okay so that is the first part of the equation right?

OpenStudy (anonymous):

Also, to actually *do* that, you're multiplying both sides by (x+23) \[A = \frac{9}{2 (x+23)} \] \[ (x+23)A = (x + 23) \left( \frac{9}{ 2(x+23)} \right) \] \[ (x+23) A = \frac{9}{2} \cancel{ \frac{(x+23)}{(x+23)}}^{ 1} \] ----------- Yeah, that's the first part of the equation :) You can divide both sides by something now to isolate x even more......

OpenStudy (anonymous):

Okay I am really really really sorry. but can you solve the equation and then after that I can see the steps because the images kind of confuse me a little bit. I am soo sorryy):

OpenStudy (anonymous):

sure :) \[ (x+23)A = \frac{9}{2} \] \[ x+23 = \frac{1}{A} \frac{9}{2} \] \[ x = \frac{9}{2A} - 23 \] follow?

OpenStudy (anonymous):

Yes thank you. Can I go ahead and write down these steps in my answer box?? :)(:

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

sure.... I don't know what that is, but the solution is correct ^^

OpenStudy (anonymous):

My answer box? Its just where I write my answer down.

OpenStudy (anonymous):

What are the answers for part b and c???